The diagonal length of a television screen is 20 in. If the width is 16 in. how high is it?
Found 2 solutions by Alan3354, ewatrrr:
Answer by Alan3354(69443)
Answer by ewatrrr(24785)


Hi
The diagonal length of a television screen is 20 in.
If the width is 16 in. how high is it?
applying the Pythagoream Theorem
h = sqrt(400-256) = sqrt(144) = 12 in, the height